# Reliable algorithm selection for machine learning-guided design

### v1

@@ Abstract
Algorithms for machine learning-guided design, or design algorithms, use machine learning-based predictions to propose novel objects with desired property values. Given a new design task—for example, to design novel proteins with high binding affinity to a therapeutic target—one must choose a design algorithm and specify any hyperparameters and predictive and/or generative models involved. How can these decisions be made such that the resulting designs are successful? This paper proposes a method for design algorithm selection, which aims to select design algorithms that will produce a distribution of design labels satisfying a user-specified success criterion—for example, that at least ten percent of designs’ labels exceed a threshold. It does so by combining designs’ predicted property values with held-out labeled data to reliably forecast characteristics of the label distributions produced by different design algorithms, building upon techniques from prediction-powered inference [2]. The method is guaranteed with high probability to return design algorithms that yield successful label distributions (or the null set if none exist), if the density ratios between the design and labeled data distributions are known. We demonstrate the method’s effectiveness in simulated protein and RNA design tasks, in settings with either known or estimated density ratios.
@@ 1 Design algorithm selection
Machine learning-guided design aims to propose novel objects, or designs, that exhibit desired values of a property of interest by consulting machine learning-based predictions of the property in place of costly and time-consuming measurements. The approach has been used to design novel enzymes that efficiently catalyze reactions of interest [18], photoreceptors with unprecedented light sensitivity for optogenetics [4], and antibodies with enhanced binding affinity to therapeutic targets [21], among many other applications. The methods used in such efforts, which we call design algorithms, are varied. Some entail sampling from a generative model that upweights objects with promising predictions [8, 7]; others start with initial candidates and iteratively introduce modifications that yield more desirable predictions [10, 46]. To propose designs, one must choose among these design algorithms. Many have consequential hyperparameter(s) that must be set, such as those that navigate a trade-off between departing from training data to achieve unprecedented predicted values and staying close enough that those predictions can be trusted [7, 21, 54]. One must also specify the predictive model that the design algorithm consults, as well as any generative model it may involve. These decisions collectively yield a design algorithm configuration, or configuration for short—for example, the design algorithm AdaLead [40] using the hyperparameter settings $r=0.2,\mu=0.02,\kappa=0.05$ and a ridge regression predictive model. The choice of the configuration dictates the resulting designs, and, consequently, whether the design effort succeeds or fails. Specifically, we say that a configuration is successful if it produces a distribution of design labels that satisfies a user-specified success criterion—for example, that at least ten percent of the design labels exceed some threshold. In this paper, we propose a solution to the problem of design algorithm selection:
To anticipate whether a configuration will be successful, one can examine predicted property values for designs that it produces. However, these predictions can be particularly error-prone, as design algorithms often produce distributions of designs shifted away from the training data in order to achieve unprecedented predicted values [14]. We propose a method for design algorithm selection that combines predictions for designs with labeled data held out from training, in a way that corrects for these errors and enables theoretical guarantees on the selected configurations. First, we formalize design algorithm selection as a multiple hypothesis testing problem. We consider a finite set of candidate design algorithm configurations, called the menu, which we describe shortly. Each configuration on the menu is affiliated with a hypothesis test of whether it satisfies the success criterion. The method then computes a statistically valid p-value for this hypothesis test, by combining held-out labeled data with predicted property values for designs generated by the configuration. Loosely speaking, it uses the labeled data to characterize how prediction error biases a p-value based on predictions alone and removes this bias, building upon techniques from prediction-powered inference [2]. Finally, the p-values are assessed with a multiple testing correction to choose a set of design algorithm configurations. The method is guaranteed with high probability to select successful configurations (or return the null set if none are found), if the ratios between the design and labeled data densities are known for all configurations on the menu. If these density ratios are unknown, we show empirically that the method still frequently selects successful configurations using estimated density ratios.
@@
The contents of the menu will depend on the task at hand. If a bespoke design algorithm has been developed specifically for the task, we may be interested in setting a key hyperparameter, such as real-valued hyperparameters that dictate how close to stay to the training data or other trusted points [36, 27, 7, 21, 16, 45]. The menu would then be a finite set of candidate values for that hyperparameter, such as a grid of values between plausible upper and lower bounds. In other cases, such as when a variety of design algorithms are appropriate for the task, we may want to consider multiple options for the design algorithm, its hyperparameter(s), and the predictive model it consults. The menu could then comprise all combinations of such options. Indeed, the menu can include or exclude any configuration we desire; we can include options for any degree of freedom whose effect on the designs we want to consider, while holding fixed those that can be reliably set by domain expertise.
@@
Our contributions are as follows. We introduce the problem of design algorithm selection, which formalizes a consequential decision faced by practitioners of machine learning-guided design and links it to the goal of producing a distribution of design labels that satisfies a criterion. We propose a method for design algorithm selection, which combines designs’ predicted property values with held-out labeled data in a principled way to reliably assess whether candidate design algorithm configurations will be successful. We provide theoretical guarantees for the configurations selected by the method, if the ratios between the design and labeled data densities are known. Finally, we demonstrate the method’s effectiveness on simulated protein and RNA design tasks, including settings where these density ratios are unknown and must be estimated.
2 Formalizing the design algorithm selection problem
@@
This section formalizes the design algorithm selection problem. The next section proposes our solution.
@@
The goal of a design task is to find novel objects in some domain, $\mathcal{X}$, whose labels in some space $\mathcal{Y}$ satisfy a desired criterion. For example, one may seek protein sequences of length $L$, $x\in\mathcal{X}=\mathcal{A}^{L}$ where $\mathcal{A}$ is the set of amino acids, whose real-valued catalytic activities for a reaction of interest, $y\in\mathcal{Y}=\mathbb{R}$, surpass some threshold. Note that it is often neither necessary nor feasible for every proposed design to satisfy the criterion—it is sufficiently useful that some of them do so. This observation guides our formalization of the success criterion, which we describe shortly. We first describe the other components of our framework.
@@
Design algorithms are methods that output novel objects whose labels are believed to satisfy the desired criterion. We focus on those that consult a predictive model, though our framework is agnostic as to how they do so; they may also use other sources of information, such as unlabeled data or biomolecular structures. An example of a design algorithm for the aforementioned enzyme design task is to sample from a generative model fit to sequences that are evolutionarily related to a known enzyme, and return the samples with desirable predictions [46].
@@ 2.1 The success criterion
Given the above components, the goal of design algorithm selection is to select a subset of configurations, $\hat{\Lambda}\subseteq\Lambda$, that satisfy the success criterion, which we now formalize.
The designs produced by any design algorithm configuration, $\lambda\in\Lambda$, are sampled from some design distribution over $\mathcal{X}$, denoted $P_{X;\lambda}$. This distribution may be specified explicitly [8, 54], or only implicitly, such as when the algorithm iteratively introduces mutations to training sequences based on the resulting predicted property values [40]. The design distribution in turn induces the design label distribution over $\mathcal{Y}$, denoted $P_{Y;\lambda}$, which is the distribution of labels of designs drawn from $P_{X;\lambda}$ and can be sampled from as follows: $x\sim P_{X;\lambda},y\sim P_{Y\mid X=x}$. Note that the labeled data and design data are related by covariate shift [38]: the distributions over $\mathcal{X}$ differ, but the conditional distribution of the label given any point is fixed.
As previously noted, the aspiration for most design efforts in practice is not that every single design satisfies a criterion, but that enough of them do so [52]. Accordingly, our framework defines success in terms of the design label distribution, $P_{Y;\lambda}$, rather than the label of any specific design. The practitioner can specify any success criterion that can be expressed as the expected value of some function of the design labels surpassing some threshold:
@@
for some $g:\mathcal{Y}\rightarrow\mathbb{R}$ and $\tau\in\mathbb{R}$ [^1] See Appendix B.1 for generalization to other success criteria; here we focus on this special case for its broad applicability. Examples of this expected function value include the mean design label when $g$ is the identity, as well as the fraction of designs whose labels exceed some value $\gamma\in\mathcal{Y}$, when $g(y)=\mathbbm{1}[y\geq\gamma]$. Using the latter, for example, a practitioner can request that at least $10\%$ of the designs’ labels exceed that of a wild-type sequence, $y_{\text{WT}}$, using the success criterion $\mathbb{E}_{Y\sim P_{Y;\lambda}}[\mathbbm{1}[Y\geq y_{\text{WT}}]]\geq 0.1$. We call a configuration successful if it yields a design label distribution that achieves the success criterion.
Our proposed method aims to select a subset of configurations from the menu such that, with guaranteed probability, every selected configuration is successful (or it returns the empty set, if no successful configuration exists on the menu). That is, for any user-specified error rate, $\alpha\in[0,1]$, the goal is to construct a subset $\hat{\Lambda}\subseteq\Lambda$ such that
@@ 3 Design algorithm selection by multiple hypothesis testing
Our method tackles design algorithm selection by approaching it as a multiple hypothesis testing problem (Alg. 1). The goal is to select a subset of configurations that are all successful, such that the error rate—the probability of incorrectly including one or more unsuccessful configurations—is at most $\alpha$. To accomplish this, for each configuration on the menu consider the null hypothesis that it is unsuccessful: $H_{\lambda}:\theta_{\lambda}:=\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]<\tau$. We compute a p-value, $p_{\lambda}$, for testing against this null hypothesis, as we describe shortly. Finally, we use the Bonferroni correction to select all configurations whose p-values are sufficiently small. This subset of configurations satisfies Eq. 2, our goal for design algorithm selection.
@@
Inputs: $N$ designs generated by each design algorithm configuration on the menu, $\{x^{\lambda}_{i}\}_{i=1}^{N},\forall\lambda\in\Lambda$; predictive models used by each configuration, $\{f_{\lambda}\}_{\lambda\in\Lambda}$; labeled data, $\{(x_{i},y_{i})\}_{i=1}^{n}$; error rate, $\alpha\in[0,1]$.
@@
The key subproblem is obtaining statistically valid p-values for testing against the null hypotheses $H_{\lambda},\lambda\in\Lambda$. These hypotheses concern expectations over the design label distributions, yet we do not have labels for the designs, only predictions. To extract information from these predictions without being misled by prediction error, we turn to prediction-powered inference [2], a framework that combines predictions with held-out labeled data to conduct valid statistical inference. Specifically, we use prediction-powered inference techniques adapted for covariate shift, due to the covariate-shift relationship between the design data and labeled data. We defer a thorough treatment to Appendix A, but conceptually, the labeled data, weighted by the density ratios between the design and labeled distributions, is used to characterize how prediction error distorts estimation of $\theta_{\lambda}$ based on predictions alone. This characterization of prediction error is then combined with the predictions for the designs to compute p-values that have either finite-sample (Alg. 3) or asymptotic (Alg. 2) validity. Using the former, the output of Algorithm 1 satisfies Eq. 2.
Theorem 1.
For any error rate, $\alpha\in[0,1]$, function $g:\mathcal{Y}\rightarrow\mathbb{R}$, and threshold $\tau\in\mathbb{R}$, Algorithm 1 using Algorithm 3 as the p-value subroutine returns a subset of configurations from the menu, $\hat{\Lambda}\subseteq\Lambda$, that satisfies
where $\theta_{\lambda}:=\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]$, and the probability is over random draws of the labeled data and the designs for all configurations on the menu.
Using asymptotically valid p-values (Alg. 2) yields an asymptotic version of the guarantee (Thm. 2). In experiments with known density ratios between the design and labeled distributions, we always achieved error rates under $\alpha$ even with asymptotically valid p-values. Since these are faster to compute, as they leverage closed-form expressions of the asymptotic null distributions, we recommend using these in practice.
@@
Inputs: Predictions for designs, $\{\hat{y}^{\prime}_{i}\}_{i=1}^{N}$; labeled data and their predictions, $\{(x_{j},y_{j},\hat{y}_{j})\}_{j=1}^{n}$.
@@
1: $\hat{\mu}_{\hat{y}^{\prime}}\leftarrow\frac{1}{N}\sum_{i=1}^{N}g(\hat{y}^{\prime}_{i})$
@@
4: $\hat{\theta}\leftarrow\hat{\mu}_{\hat{y}^{\prime}}+\hat{\Delta}$
5: $\hat{\sigma}^{2}_{\hat{y}^{\prime}}\leftarrow\frac{1}{N}\sum_{i=1}^{N}(g(\hat{y}^{\prime}_{i})-\hat{\mu}_{\hat{y}^{\prime}})^{2}$
6: $\hat{\sigma}^{2}_{\hat{y}-y}\leftarrow\frac{1}{n}\sum_{j=1}^{n}\left(w_{j}[g(y_{j})-g(\hat{y}_{j})]-\hat{\Delta}\right)^{2}$
7: $P\leftarrow 1-\Phi\left(\left(\hat{\theta}-\tau\right)\left/\right.\sqrt{\frac{\hat{\sigma}^{2}_{\hat{y}^{\prime}}}{N}+\frac{\hat{\sigma}^{2}_{\hat{y}-y}}{n}}\right)$triangleright $\Phi$ is the standard normal CDF
@@
Note that an error rate of zero can be trivially achieved by returning the empty set—that is, by never selecting any configuration. This is a legitimate outcome if there are no successful configurations on the menu, but otherwise, we will show empirically that our method also exhibits a high selection rate of returning nonempty sets. On the other hand, if our method selects multiple configurations then one can safely use any (or any mixture) of them with the same high-probability guarantee of achieving the success criterion. In particular, one can further narrow down $\hat{\Lambda}$ using additional criteria—for example, picking the selected configuration that produces the most diverse designs, in order to hedge against unknown future desiderata.
@@
Computing the p-values requires the density ratio between the design and labeled distributions, $p_{X;\lambda}(x_{i})/p_{\text{lab}}(x_{i})$, for every configuration and every labeled instance. Important settings in biological sequence design where the design density can be evaluated are when designs are sampled from autoregressive generative models [39] or when the design distribution is a product of independent categorical distributions per sequence site, as in library design [51, 54]. Labeled sequence data is often generated by adding random substitutions to wild type sequences [7, 10] or by recombining segments of several “parent” sequences [33, 4], in which case $p_{\text{lab}}$ is also explicitly known. Valid p-values can also be computed if both $p_{X;\lambda}$ and $p_{\text{lab}}$ are only known up to normalizing factors, such as when sequences are generated by Potts models [34, 16] (Appendix B.2). In other settings, however, these density ratios need to be approximated. In experiments with unknown density ratios, we use multinomial logistic regression-based density ratio estimation [44] and show that our method still empirically outperforms others in selecting successful configurations. Although Theorems 1 or 2 no longer apply in this setting, they inform us what guarantees we can recover, to the extent that the density ratios are approximated well.
What does it mean to go “too far?”
For any given configuration, $\lambda\in\Lambda$, the further apart the design and labeled distributions are, the higher the variance of the density ratios that are used to weight the labeled data. This reduces the effective sample size of the labeled data in characterizing prediction error, which leads to higher uncertainty about the value of $\theta_{\lambda}$. Consequently, even if the configuration is successful, it may not be selected if the design and labeled distributions are too far apart. Indeed, this is desirable behavior: the method essentially identifies where over $\mathcal{X}$ we lack sufficient statistical evidence of achieving the success criterion. It returns the empty set if the predictions and labeled data do not collectively provide adequate evidence that any configuration on the menu is successful.

### v2

@@ Abstract
Algorithms for machine learning-guided design, or design algorithms, use machine learning-based predictions to propose novel objects with desired property values. Given a new design task—for example, to design novel proteins with high binding affinity to a therapeutic target—one must choose a design algorithm and specify any hyperparameters and predictive and/or generative models involved. How can these decisions be made such that the resulting designs are successful? This paper proposes a method for design algorithm selection, which aims to select design algorithms that will produce a distribution of design labels satisfying a user-specified success criterion—for example, that at least ten percent of designs’ labels exceed a threshold. It does so by combining designs’ predicted property values with held-out labeled data to reliably forecast characteristics of the label distributions produced by different design algorithms, building upon techniques from prediction-powered inference (Angelopoulos et al.,, 2023). The method is guaranteed with high probability to return design algorithms that yield successful label distributions (or the null set if none exist), if the density ratios between the design and labeled data distributions are known. We demonstrate the method’s effectiveness in simulated protein and RNA design tasks, in settings with either known or estimated density ratios.
@@ 1 Design Algorithm Selection
Machine learning-guided design aims to propose novel objects, or designs, that exhibit desired values of a property of interest by consulting machine learning-based predictions of the property in place of costly and time-consuming measurements. The approach has been used to design novel enzymes that efficiently catalyze reactions of interest (Greenhalgh et al.,, 2021), photoreceptors with unprecedented light sensitivity for optogenetics (Bedbrook et al.,, 2019), and antibodies with enhanced binding affinity to therapeutic targets (Gruver et al.,, 2023), among many other applications. The methods used in such efforts, which we call design algorithms, are varied. Some entail sampling from a generative model that upweights objects with promising predictions (Brookes et al.,, 2019; Biswas et al.,, 2021); others start with initial candidates and iteratively introduce modifications that yield more desirable predictions (Bryant et al.,, 2021; Thomas et al.,, 2025). To propose designs, one must choose among these design algorithms. Many have consequential hyperparameter(s) that must be set, such as those that navigate a trade-off between departing from training data to achieve unprecedented predicted values and staying close enough that those predictions can be trusted (Biswas et al.,, 2021; Gruver et al.,, 2023; Zhu et al.,, 2024). One must also specify the predictive model that the design algorithm consults, as well as any generative model it may involve. These decisions collectively yield a design algorithm configuration, or configuration for short—for example, the design algorithm AdaLead (Sinai et al.,, 2020) using the hyperparameter settings $r=0.2,\mu=0.02,\kappa=0.05$ and a ridge regression predictive model. The choice of the configuration dictates the resulting designs, and, consequently, whether the design effort succeeds or fails. Specifically, we say that a configuration is successful if it produces a distribution of design labels that satisfies a user-specified success criterion—for example, that at least ten percent of the design labels exceed some threshold. In this paper, we propose a solution to the problem of design algorithm selection:
To anticipate whether a configuration will be successful, one can examine predicted property values for designs that it produces. However, these predictions can be particularly error-prone, as design algorithms often produce distributions of designs shifted away from the training data in order to achieve unprecedented predicted values (Fannjiang and Listgarten,, 2024). We propose a method for design algorithm selection that combines predictions for designs with labeled data held out from training, in a way that corrects for these errors and enables theoretical guarantees on the selected configurations. First, we formalize design algorithm selection as a multiple hypothesis testing problem. We consider a finite set of candidate design algorithm configurations, called the menu, described shortly. Each configuration is affiliated with a hypothesis test of whether it satisfies the success criterion. The method then computes a statistically valid p-value for this hypothesis test, by combining held-out labeled data with predicted property values for designs generated by the configuration. Loosely speaking, it uses the labeled data to characterize how prediction error biases a p-value based on predictions alone and removes this bias, building upon techniques from prediction-powered inference (Angelopoulos et al.,, 2023). Finally, the p-values are assessed with a multiple testing correction to choose a set of design algorithm configurations. The method is guaranteed with high probability to select successful configurations (or return the null set if there are none), if the ratios between the design and labeled data densities are known for all configurations on the menu. If these density ratios are unknown, we show empirically that the method still frequently selects successful configurations using estimated density ratios.
@@
The contents of the menu will depend on the task at hand. If a bespoke design algorithm has been developed specifically for the task, we may be interested in setting a key hyperparameter, such as real-valued hyperparameters that dictate how close to stay to the training data or other trusted points (Schubert et al.,, 2018; Linder et al.,, 2020; Biswas et al.,, 2021; Gruver et al.,, 2023; Fram et al.,, 2024; Tagasovska et al.,, 2024). The menu would then be a finite set of candidate values for that hyperparameter, such as a grid of values between plausible upper and lower bounds. In other cases, such as when a variety of design algorithms are appropriate for the task, we may want to consider multiple options for the design algorithm, its hyperparameter(s), and the predictive model it consults. The menu could then comprise all combinations of such options. Indeed, the menu can include or exclude any configuration we desire; we can include options for any degree of freedom whose effect on the designs we want to consider, while holding fixed those that can be reliably set by domain expertise.
@@
Our contributions are as follows. We introduce the problem of design algorithm selection, which formalizes a consequential decision faced by practitioners of machine learning-guided design and connects it to the goal of producing a distribution of design labels that satisfies a criterion. We propose a method for design algorithm selection, which combines designs’ predicted property values with held-out labeled data in a principled way to reliably assess whether candidate design algorithm configurations will be successful. We provide theoretical guarantees for the configurations selected by the method, if the ratios between the design and labeled data densities are known. Finally, we show the method’s effectiveness in simulated protein and RNA design tasks, including settings where these density ratios are unknown and must be estimated.
2 Problem Formalization
@@
This section formalizes the design algorithm selection problem, and the next proposes our solution.
@@
The goal of a design task is to find novel objects in some domain, $\mathcal{X}$, whose labels in some space $\mathcal{Y}$ satisfy a desired criterion. For example, one may seek length-$L$ protein sequences, $x\in\mathcal{X}=\mathcal{A}^{L}$ where $\mathcal{A}$ is the set of amino acids, whose real-valued catalytic activities for a reaction of interest, $y\in\mathcal{Y}=\mathbb{R}$, surpass some threshold. Note that it is often neither necessary nor feasible for every proposed design to satisfy the criterion—it is sufficiently useful that some of them do (Wheelock et al.,, 2022). This observation guides our formalization of the success criterion, which we describe shortly. We first describe the other components of our framework.
@@
Design algorithms are methods that output novel objects whose labels are believed to satisfy the desired criterion. We focus on those that consult a predictive model, though our framework is agnostic as to how they do so; they may also use other sources of information, such as unlabeled data or biomolecular structures. An example of a design algorithm for the aforementioned enzyme design task is to sample from a generative model fit to sequences that are evolutionarily related to a known enzyme, and return the samples with desirable predicted catalytic activities (Thomas et al.,, 2025).
@@ 2.1 The Success Criterion
Given the above components, the goal of design algorithm selection is to select a subset of configurations, $\hat{\Lambda}\subseteq\Lambda$, that satisfy the success criterion, which we now formalize. The designs produced by any design algorithm configuration, $\lambda\in\Lambda$, are sampled from some design distribution over $\mathcal{X}$, denoted $P_{X;\lambda}$. This distribution may be specified explicitly (Brookes et al.,, 2019; Zhu et al.,, 2024), or only implicitly, such as when the algorithm iteratively introduces mutations to training sequences based on the resulting predicted property values (Sinai et al.,, 2020). The design distribution in turn induces the design label distribution over $\mathcal{Y}$, denoted $P_{Y;\lambda}$, which is the distribution of design labels and can be sampled from as follows: $x\sim P_{X;\lambda},y\sim P_{Y\mid X=x}$. Note that the labeled data and design data are related by covariate shift (Shimodaira,, 2000): the distributions over $\mathcal{X}$, $P_{\text{lab}}$ and $P_{X;\lambda}$, differ, but the conditional distribution of the label for any point, $P_{Y\mid X=x}$ for any $x$, is fixed.
As previously noted, the aspiration for most design endeavors in practice is not that every single design satisfies a criterion, but that enough of them do so (Wheelock et al.,, 2022). Accordingly, our framework defines success in terms of the design label distribution, $P_{Y;\lambda}$, rather than the label of any specific design. The practitioner can specify any success criterion that requires the expected value of some function of the design labels to surpass some threshold:
@@
for some $g:\mathcal{Y}\rightarrow\mathbb{R}$ and $\tau\in\mathbb{R}$ [^1] See Appendix B.1 for generalization to other success criteria; here we focus on this special case for its broad applicability. We call $\theta_{\lambda}$ the population-level metric. Examples include the mean design label when $g$ is the identity, as well as the fraction of designs whose labels exceed some value $\gamma\in\mathcal{Y}$, when $g(y)=\mathbbm{1}[y\geq\gamma]$. For example, a practitioner can request that at least ten percent of the designs’ labels exceed that of a wild type, $y_{\text{WT}}$, using the success criterion $\mathbb{E}_{Y\sim P_{Y;\lambda}}[\mathbbm{1}[Y\geq y_{\text{WT}}]]\geq 0.1$. We call a configuration successful if it yields a design label distribution that achieves the success criterion.
Our goal is to select a subset of configurations from the menu such that, with guaranteed probability, every selected configuration is successful (or to return the empty set, if no successful configuration exists on the menu). That is, for any user-specified error rate $\alpha\in[0,1]$, we aim to construct a subset $\hat{\Lambda}\subseteq\Lambda$ such that
@@ 3 Design Algorithm Selection by Multiple Hypothesis Testing
Our method approaches design algorithm selection as a multiple hypothesis testing problem (Alg. 1). The goal is to select a subset of configurations that are all successful, such that the error rate—the probability of incorrectly including one or more unsuccessful configurations—is at most $\alpha$. To accomplish this, for each configuration on the menu, consider the null hypothesis that it is unsuccessful: $H_{\lambda}:\theta_{\lambda}:=\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]<\tau$. We compute a p-value, $p_{\lambda}$, for testing against this null hypothesis, as we describe shortly. Finally, we use the Bonferroni correction to select all configurations with sufficiently small p-values. This subset of configurations satisfies Eq. 2, our goal for design algorithm selection.
The key subproblem is obtaining statistically valid p-values for testing against the null hypotheses $H_{\lambda},\lambda\in\Lambda$. These hypotheses concern the design label distributions, yet we do not have labels for the designs, only predictions. To extract information from these predictions without being misled by prediction error, we turn to prediction-powered inference (Angelopoulos et al.,, 2023), a framework that combines predictions with held-out labeled data to conduct valid statistical inference. Specifically, we use prediction-powered inference techniques adapted for covariate shift, due to the covariate-shift relationship between the design data and labeled data. We defer a thorough treatment to Appendix A.1, but conceptually, the labeled data, weighted by the density ratios between the design and labeled distributions, is used to characterize how prediction error distorts estimation of the population-level metric, $\theta_{\lambda}$, based on predictions alone. This error characterization is then combined with the predictions for the designs to compute p-values that have either asymptotic (Alg. 2) or finite-sample (Alg. 3) validity. Using the latter, the output of Algorithm 1 satisfies Eq. 2.
@@
Inputs: $N$ designs generated by each design algorithm configuration on the menu, $\{x^{\lambda}_{i}\}_{i=1}^{N},\forall\lambda\in\Lambda$; predictive models used by each configuration, $\{f_{\lambda}\}_{\lambda\in\Lambda}$; labeled data, $\{(x_{j},y_{j})\}_{j=1}^{n}$; error rate, $\alpha\in[0,1]$.
@@
Inputs: Predictions for designs, $\{\hat{y}^{\lambda}_{i}\}_{i=1}^{N}$; labeled data and their predictions, $\{(x_{j},y_{j},\hat{y}_{j})\}_{j=1}^{n}$.
@@
1: $\hat{\mu}\leftarrow\frac{1}{N}\sum_{i=1}^{N}g(\hat{y}^{\lambda}_{i})$
@@
4: $\hat{\theta}\leftarrow\hat{\mu}+\hat{\Delta}$
5: $\hat{\sigma}^{2}_{\text{pred}}\leftarrow\frac{1}{N}\sum_{i=1}^{N}(g(\hat{y}^{\lambda}_{i})-\hat{\mu})^{2}$
6: $\hat{\sigma}^{2}_{\text{err}}\leftarrow\frac{1}{n}\sum_{j=1}^{n}\left(w_{j}[g(y_{j})-g(\hat{y}_{j})]-\hat{\Delta}\right)^{2}$
7: $P\leftarrow 1-\Phi\left(\left(\hat{\theta}-\tau\right)\left/\right.\sqrt{\frac{\hat{\sigma}^{2}_{\text{pred}}}{N}+\frac{\hat{\sigma}^{2}_{\text{err}}}{n}}\right)$triangleright $\Phi$ is the standard normal CDF
Theorem 3.1.
For any error rate $\alpha\in[0,1]$, function $g:\mathcal{Y}\rightarrow\mathbb{R}$, and threshold $\tau\in\mathbb{R}$, Algorithm 1 using Algorithm 3 as the p-value subroutine returns a subset of configurations from the menu, $\hat{\Lambda}\subseteq\Lambda$, that satisfies
where $\theta_{\lambda}:=\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]$, and the probability is over random draws of labeled data and designs from all configurations on the menu.
Using asymptotically valid p-values (Alg. 2) yields an asymptotic version of the guarantee (Thm. A.1). In experiments with known density ratios between the design and labeled distributions, we always achieved error rates under $\alpha$ even with asymptotically valid p-values. Since these are faster to compute, as they leverage closed-form representations of the asymptotic null distributions, we recommend using these in practice.
@@
Note that an error rate of zero can be trivially achieved by returning the empty set—that is, by not selecting any configuration. This is a legitimate outcome if there are no successful configurations on the menu, but otherwise, we will show empirically that our method also exhibits a high selection rate, or rate of returning nonempty sets. On the other hand, if our method selects multiple configurations then one can safely use any (or any mixture) of them with the same high-probability guarantee of success. In particular, one can further narrow down $\hat{\Lambda}$ using additional criteria—for example, picking the selected configuration that produces the most diverse designs, in order to hedge against unknown future desiderata.
@@
Computing the p-values requires the density ratio between the design and labeled distributions, $p_{X;\lambda}(x_{i})/p_{\text{lab}}(x_{i})$, for every configuration and every labeled instance. Important settings in biological sequence design where the design density can be evaluated are when designs are sampled from autoregressive generative models (Shin et al.,, 2021) or when the design distribution is a product of independent categorical distributions per sequence site (Weinstein et al.,, 2022; Zhu et al.,, 2024). Labeled sequence data is often generated by adding random substitutions to wild-type sequences (Biswas et al.,, 2021; Bryant et al.,, 2021) or by recombining segments of several “parent” sequences (Romero et al.,, 2013; Bedbrook et al.,, 2019), in which case $p_{\text{lab}}$ is also explicitly known. Valid p-values can also be computed if both $p_{X;\lambda}$ and $p_{\text{lab}}$ are only known up to normalizing factors, such as when sequences are generated by Potts models (Russ et al.,, 2020; Fram et al.,, 2024) (Appendix B.2). In other settings, however, these density ratios need to be approximated. In experiments with unknown density ratios, we use multinomial logistic regression-based density ratio estimation (Srivastava et al.,, 2023) and show that our method still empirically outperforms others in selecting successful configurations. Although Theorems 3.1 or A.1 no longer apply in this setting, they inform us what guarantees we can recover, to the extent that the density ratios are approximated well.
When have we gone “too far?”
For each configuration, the further apart the design and labeled distributions are, the higher the variance of the density ratios that are used to weight the labeled data. This reduces the effective sample size of the labeled data in characterizing prediction error, which leads to higher uncertainty about the value of the population-level metric, $\theta_{\lambda}$. Consequently, even if the configuration is successful, it may not be selected if the design and labeled distributions are too far apart. Indeed, this is desirable behavior: the method essentially identifies where over $\mathcal{X}$ we lack sufficient statistical evidence of achieving the success criterion. It returns the empty set if the predictions and labeled data do not collectively provide adequate evidence that any configuration on the menu is successful.
Another factor to how many configurations are selected is the multiple testing correction, as it determines what counts as adequate evidence of success. A conceptual strength of framing design algorithm selection as a multiple testing problem is that any multiple testing procedure that controls family-wise error rate (FWER) can be used in Line 6 of Algorithm 1. Our instantiation uses the Bonferroni correction for full generality, as it does not require any assumptions about how different configurations are related, and it yielded reasonably high selection rates with menus of a few hundred configurations in experiments. However, one could also substitute FWER-controlling procedures that account for hierarchical (Bretz et al.,, 2009) or correlation structure (Dudoit et al.,, 2003) in the menu, which could yield less conservative multiplicity corrections and therefore higher selection rates.